OpenStudy (anonymous):

The Empire State Building reaches one thousand two hundred fifty feet at the top floor. Keira has a model of this landmark building that is 8 inches tall. What is the scale between the model and the actual building?

OpenStudy (anonymous):

My options are 1inch:8feet 1inch:156.25 1foot:8inches 1foot:156.25inches

OpenStudy (anonymous):

@Jesuslover14

OpenStudy (anonymous):

U HELP ME ILL HELP U 11. (03.06 LC) Choose the equation below that represents the line that passes through the point (-7, 2) and has a slope of −3. (5 points) y − 2 = −3(x + 7) y + 7 = −3(x − 2) y + 2 = −3(x − 7) y − 7 = −3(x + 2) Score: 0 of 5

OpenStudy (anonymous):

I BELIEVE IT IS THE THE THIRD OPTION

OpenStudy (anonymous):

COULD U HELP ME NOW

OpenStudy (anonymous):

i dont know sorry

Directrix (directrix):

@Jesuslover14 Please create your own threads in which you post your questions.